OpenStudy (anonymous):

The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. A = 2/3 ft.; c = 4 ft. A = sq. ft.

OpenStudy (anonymous):



OpenStudy (pfenn1):

Area of a triangle \[A=\frac12 bh\]

OpenStudy (anonymous):

area = bh/2 b is the base length h is the height here a is your base and b is your height so you should get (1/2)(2/3)(4) =4/3

OpenStudy (pfenn1):

@cerezas, where did you get the value for b?

OpenStudy (anonymous):

misread c>.< but you can use pythagoras to find it

OpenStudy (anonymous):

4^2 - (4/3)^2 - b^2 16 - 16/9 = b^2 b^2 = 128/9 b = 8(sqrt(2))/9